HUBBLE SPACE TELESCOPE DAILY REPORT #1854 PERIOD COVERED: 0000Z (UTC) 04/10/97 - 0000Z (UTC) 04/11/97 Daily Status Report as of 101/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: ************************************* Zero gyro safemode occurred at approximately 099/171712Z. The = following observations could not be executed due to the safemode event: ************************************** 1.1 NIC 1/2/3 7049 (NICMOS Absolute Photometry Test) 1.2 NIC 1/2 7128 (Early Release Observations of Allen's Source in NGC-2264) 1.3 NIC /1/2 7120 (IRC+10216: The Second Brightest Non- Solar System Object in the IR Sky) 1.4 Two Sets of WF/PC-2 6346 (Snapshot Survey of the Ultraluminous IRAS Galaxy Sample-Part 1 High) 1.5 WF/PC-2 6363 (The Environments and Host Galaxies of BL-Lac Objects) 1.6 WF/PC-2 6361 (High Resolution Imaging of X-Ray Selected AGNs) 1.7 WF/PC-2 6359 (Core Properties of the Bulges of Spiral Galaxies) 1.8 WF/PC-2 6578 (Optical Morphology of NGC 5128: X-Ray Ridges and the Geometry of the Dust Lane) 1.9 STIS/CCD 7092 (CCD Dark Rate and Read Noise) 1.10 WF/PC-2 6887 (A Photometric Search for Faint Companions to Nearby Stars) 1.11 STIS/CCD 7130 (STIS/Spectroscopy of the Brown Dwarf GL229B) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 6 Successful : 0 Scheduled Reacquisitions : 3 Successful : 0 There were no acquisitions or reacquisitions due to the safing = event. 2.2 FHST Updates: Scheduled : 3 Successful : 0 No updates were performed due to the safing event. 2.3 Operations Notes: PSEA configuration was modified to select RGA 1 at 100/1448Z. STIS was recovered from safemode at 100/1625Z. The health and safety load needed for spacecraft recovery was = initiated at 100/2221Z. The recovery to normal mode was completed at 100/2300Z. SSR memory monitors were reset at 101/0556Z. The SSR row error = test was conducted at 101/0853Z. An SSR software retry patch was = partially installed at 101/0916Z and at 101/1004Z. The remainder of the = patch procedure will be completed later. After the NICMOS thermal control setpoint was corrupted in the SSA = (HSTAR 6022), it was reset at 101/1448Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Resumption of normal science and SMOV operations is scheduled for = approximately 101/1600z (Saturday, April 12,1997). /DMH